OpenStudy (anonymous):

Calculate the mass of barium chlorate formed from .100 liter of 3.00 M chloric acid solution plus excess barium hydroxide.

OpenStudy (aaronq):

1) Write the equation for the reaction 2) Convert hydrochloric acid solution to moles (use \(Molarity=\dfrac{n_{solute}}{L_{solution}}\)) 3) Use a ratio of the moles and the coefficients \(\sf \dfrac{n_{HCl}}{HCl's~coefficient}=\dfrac{n_{BaCl_2}}{BaCl_2's~coefficient}\) solve for moles of barium chlorate. 4) Convert the moles to mass

OpenStudy (anonymous):

Isnt barium chlorate Ba(ClO3)2?

OpenStudy (anonymous):

Hey

OpenStudy (anonymous):

Hoe would i calculate that? By doing the same process?

OpenStudy (anonymous):

How

OpenStudy (aaronq):

oh sorry, i messed up the formula. but yes, it's the same process regardless of what's reacting.